### v1

# Casper: Causality-Aware Spatiotemporal Graph Neural Networks for Spatiotemporal Time Series Imputation

Abstract.
Spatiotemporal time series is the foundation of understanding human activities and their impacts, which is usually collected via monitoring sensors placed at different locations. The collected data usually contains missing values due to various failures, e.g., mechanical damage, which have significant impact on data analysis. To impute the missing values, a lot of methods have been introduced in the literature. When recovering a specific data point, most existing methods tend to take into consideration all the information relevant to that point regardless of whether they have a cause-and-effect relationship. During data collection, it is inevitable that some unknown confounders are included, e.g., background noise in time series and non-causal shortcut edges in the constructed sensor network. These confounders could open backdoor paths between the input and output – in other words, they establish non-causal correlations between the input and output. Over-exploiting these non-causal correlations could result in overfitting and make the model vulnerable to noises. In this paper, we first revisit spatiotemporal time series imputation from a causal perspective, which shows the causal relationships among the input, output, embeddings and confounders. Next, we show how to block the confounders via the frontdoor adjustment. Based on the results of the frontdoor adjustment, we introduce a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper), which contains a novel Spatiotemporal Causal Attention (SCA) and a Prompt Based Decoder (PBD). PBD could reduce the impact of confounders and SCA could discover the sparse causal relationships among embeddings. Theoretical analysis reveals that SCA discovers causal relationships based on the values of gradients. We evaluate Casper on three real-world datasets, and the experimental results show that Casper could outperforms the baselines and could effectively discover the causal relationships.
Spatiotemporal Time Series Imputation, Spatiotemporal Graph Neural Network, Causal Attention
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Spatiotemporal time series is the cornerstone of analyzing and understanding the patterns and impact of human activities, e.g., traffic flows (Li et al., 2018) and greenhouse gas emissions (Zheng et al., 2015), which is the basis for formulating public policies, e.g., the traffic transportation policy and the clean air act. Spatiotemporal time series is usually collected by monitoring sensors. Traffic flow time series, e.g., speed and occupancy, is recorded by the radar sensors placed along roads. Air pollution time series, e.g., concentrations of CO and PM2.5, is collected from air quality monitoring sites across cites.
In the real world, it is not uncommon that the collected spatiotemporal time series is incomplete with a lot of missing data due to various failures, e.g., the Internet is disconnected or the sensors have mechanical damages. The missing data usually has significant impact on the process and conclusion of data analysis. Therefore, how to reconstruct the missing data from the observed data, i.e., imputation, is a fundamental problem of spatiotemporal time series analysis. In recent years, deep learning methods become the mainstream for time series imputation. Most existing deep time series imputation methods (Cao et al., 2018; Luo et al., 2018, 2019) use Recurrent Neural Network (RNN) to capture the temporal dynamics of time series and autoregressively recover the missing data by the predicted values. Recent deep learning methods (Yıldız et al., 2022) propose to use non-autoregressive structures, e.g., Transformer (Vaswani et al., 2017), to avoid the progressive error propagation incurred via the autoregression in RNN by concurrently considering the entire input context. However, these methods only consider the temporal patterns yet overlook the spatial relationships among sensors, e.g., geographical distances. To further account for spatial relationships, spatiotemporal graph neural networks are introduced (Jing et al., 2021; Huang et al., 2020; Marisca et al., 2022). Although these methods have achieved impressive performance for recovering the missing values, when recovering a data point, they tend to include all the available information related to the data point as reference without distinguishing whether there is a causal relationship between them.
When collecting datasets, it is inevitable that some unknown confounders (Pearl and Mackenzie, 2018) are included. For example, the background noise might be recorded, and non-causal shortcut edges might be built for two sensors. Let’s take the air monitoring sensor network as a concrete example to understand what are non-causal edges. A common practice to build the network is adding an edge for two sensors if their distance is below a threshold (Marisca et al., 2022). Although simple and usually effective, the distance based network does not necessarily imply the real causality between sensors. In the real world, air flow between two locations could be influenced by other factors, e.g., wind direction and terrain. An example is shown in Figure 6 in the Experiments section. Simply exploiting the shortcut edges without discovering the causality, could make the model overfit the training data and be vulnerable to the noise during the test.
To reduce the negative effects brought by the confounders, we first review the process of spatiotemporal time series imputation from a causal perspective and show the causal relationships among the input, output, embeddings and confounders. As shown in Figure 1, the confounders could open shortcut backdoor paths between the input and output. Then, we show how to eliminate the backdoor paths opened by confounders via the frontdoor adjustment (Pearl and Mackenzie, 2018). Based on the results of the frontdoor adjustment, we introduce a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper), which is equipped with a novel Spatiotemporal Causal Attention (SCA) and a Prompt Based Decoder (PBD). The proposed PBD effectively reduces the impact of the unknown confounders by injecting the global context information of datasets to the embeddings. Inspired by (Jia et al., 2022), which uses learnable prompts to capture the contextual information of downstream tasks when tuning visual models, PBD leverages prompts to learn the contextual information of datasets automatically rather than employing external models to approximate the context. To further enforce sparse causality between embeddings, we introduce SCA, which determines the cause-and-effect relationship via a causal gate. It can be theoretically proven that the proposed causal gate (1) enforces the sparsity since it converges to 0 or 1; (2) is actually a gradient-based explanation similar to (Selvaraju et al., 2017), which determines the causality based on the values of gradients. We extensively evaluate Casper on three real-world spatiotemporal time series imputation datasets. The experimental results show that Casper could significantly outperform baselines and could effectively discover the causality.
The major contributions are summarized as follows:
  • We review the spatiotemporal time series imputation task from a causal perspective, where we show the problems of the undesired confounders. Then, we show how to eliminate confounders via the frontdoor adjustment.
  • We propose a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper) based on the frontdoor adjustment. Casper is equipped with a novel Spatiotemporal Causal Attention (SCA) and a Prompt Based Decoder (PBD). PBD effectively blocks the backdoor paths and SCA explicitly reveals the causality between emebeddings.
  • We provide theoretical analysis to deeply understand how Casper determines causal and non-causal relationships.
  • We evaluate Casper on three real-world datasets. The experimental results show that Casper could outperform the baselines and effectively discover causal relationships.
2. Preliminary
In this section, we briefly introduce the definitions of spatiotemporal time series and spatiotemporal time series imputation. We also review the definitions of Granger causality and attention function.
Definition 0 (Incomplete Spatiotemporal Time Series).
We denote an incomplete spatiotemporal time series with missing values as $\mathbf{G}=(\mathbf{X},\mathbf{A},\mathbf{M})$, where $\mathbf{X}\in\mathbb{R}^{N\times T}$ is the multivariate time series collected from $N$ sensors with totally $T$ steps, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix of the sensor network, $\mathbf{M}\in\{0,1\}^{N\times T}$ is the binary mask and 0/1 denotes the absence/presence of a data point.
Definition 0 (Spatiotemporal Time Series Imputation).
Given an incomplete spatiotemporal time series $\mathbf{G}=(\mathbf{X},\mathbf{A},\mathbf{M})$, we denote $\mathbf{Y}\in\mathbf{R}^{N\times T}$ as the complete time series of $\mathbf{X}$, such that $\mathbf{X}=\mathbf{M}\odot{\mathbf{Y}}$, where $\odot$ is the Hadamard product. The task is to build a function $\hat{\mathbf{Y}}=f(\mathbf{G})$ to minimize the reconstruction error, e.g., Mean Absolute Error (MAE), between $\hat{\mathbf{Y}}$ and $\mathbf{Y}$.
Definition 0 (Granger Causality (Granger, 1969; Cheng et al., 2022)).
Let $\mathbf{X}\in\mathbb{R}^{N\times T}$ be the values of past $T$ steps of $N$ time series, and $\hat{y}_{i,T+1}=f_{i}(\mathbf{X})\in\mathbb{R}$ be the time series forecasting function predicting the future value of the $i$-th time series at the $T+1$ step, where $i\in[1,\dots,N]$. The $j$-th time series is said to Granger cause the $i$-th time series if there exists a point $x^{\prime}_{j,t}\neq x_{j,t}$, $t\in[1,\dots,T]$, such that $f_{i}(\mathbf{X}^{\prime})\neq f_{i}(\mathbf{X})$, where $\mathbf{X}^{\prime}$ is obtained by replacing $x_{j,t}$ in $\mathbf{X}$ with $x^{\prime}_{j,t}$.
Generally speaking, if $x_{j,t}$ has impact on the prediction of the future value of the $i$-th time series, then the $j$-th time series Granger causes the $i$-th time series. In the case that $f_{i}$ is a linear model:
(1) $\hat{y}_{i,T+1}=f_{i}(\mathbf{X})=\sum_{i^{\prime},t^{\prime}}\alpha_{i^{\prime},t^{\prime}}x_{i^{\prime},t^{\prime}}.$
If the coefficient $\alpha_{j,t}\neq 0$, then the $j$-th time series Granger causes the $i$-th time series.
Definition 0 (Attention Function).
Let $\mathbf{q}$ and $\{\mathbf{k}_{i}\}_{i=1}^{N}$ be the $d$-dimensional query and key embeddings, let $s$ and $f_{v}$ be the scoring and message functions, then the attention function is defined as:
(2) $\mathbf{h}=\sum_{i=1}^{N}\alpha_{i}\mathbf{v}_{i},\quad\alpha_{i}=\frac{\exp(s(\mathbf{q},\mathbf{k}_{i}))}{\sum_{i^{\prime}=1}^{N}\exp(s(\mathbf{q},\mathbf{k}_{i^{\prime}}))},\quad\mathbf{v}_{i}=f_{v}(\mathbf{q},\mathbf{k}_{i}),$
where $\mathbf{h}\in\mathbb{R}^{d}$ is the output, and $\mathbf{v}_{i}$ is the message from $\mathbf{k}_{i}$ to $\mathbf{q}$.
3. Methodology
In this section, we first provide a causal view of the spatiotemporal imputation task, and show how to eliminate the unknown confounds by frontdoor adjustment, based on which, we introduce a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper). Finally, we provide further analysis of Casper.
3.1. Causal View of Spatiotemporal Imputation
Given an incomplete spatiotemporal time series $\mathbf{G}=(\mathbf{X},\mathbf{A},\mathbf{M})$, a standard deep imputation model $f=f^{D}\circ f^{E}$, where $f^{E}$, $f^{D}$ are the encoder and decoder, works as follows: 1) $f^{E}$ extracts embeddings =$\mathcal{H}=\{\mathbf{h}_{i,t}\}_{i=1,t=1}^{N,T}$ from $\mathbf{G}$, 2) $f^{D}$ generates predictions $\{\hat{y}_{i,t}\}_{i=1,t=1}^{N,T}$ based on $\mathcal{H}$ to recover $\mathcal{Y}=\{{y}_{i,t}\}_{i=1,t=1}^{N,T}$. The model $f$ is trained by a reconstruction error e.g., MAE or RMSE. Since minimizing MAE (or RMSE) is equivalent to maximizing the log-likelihood of Laplace (or Gaussian) distribution (Hodson, 2022), and thus we can view the objective of spatiotemporal imputation as maximizing $P(\mathcal{Y}|\mathcal{G})$. Most existing studies focus on maximizing $P(\mathcal{Y}|\mathcal{G})$ yet few discuss the cause-and-effect relationship between $\mathcal{G}$ and $\mathcal{Y}$. In this paper, we study their causality based on the Structure Causal Model (SCM) (Pearl and Mackenzie, 2018).
Figure 1. The Structure Causal Model (SCM) for spatiotemporal imputation. The frontdoor adjustment removes the edge between confounder $\mathcal{C}$ and input $\mathcal{G}$.
3.1.1. Structure Causal Model
During data collection, it is inevitable that some unknown confounders $\mathcal{C}$ are included in the datasets, which influence both $\mathcal{G}$ and $\mathcal{Y}$. For example, sensors might record random background noise, and the constructed sensor network might contain shortcut edges. The undesired information might bridge the input $\mathcal{G}_{M}$ and the output $\mathcal{Y}$ with spurious correlations, which could lead to overfitting and make the model error-prone.
The causal relationship between $\mathcal{G}$, $\mathcal{H}$, $\mathcal{Y}$ and $\mathcal{C}$ can be modeled by the Structure Causal Model (SCM) (Pearl and Mackenzie, 2018) shown in Figure 1. First, it is evident that $\mathcal{C}$ and $\mathcal{H}$ are not d-separable (Murphy, 2012), since $\mathcal{C}$ can reach $\mathcal{H}$ via the path $\mathcal{C}\rightarrow\mathcal{G}\rightarrow\mathcal{H}$. This means that $\mathcal{H}$ and $\mathcal{C}$ are not independent and thus $\mathcal{H}$ contains information of $\mathcal{C}$. Second, besides $\mathcal{G}\rightarrow\mathcal{H}\rightarrow\mathcal{Y}$, $\mathcal{C}$ introduces backdoor paths between $\mathcal{G}$ and $\mathcal{Y}$, as well as $\mathcal{H}$ and $\mathcal{Y}$: $\mathcal{G}\leftarrow\mathcal{C}\rightarrow\mathcal{Y}$, $\mathcal{H}\leftarrow\mathcal{G}\leftarrow\mathcal{C}\rightarrow\mathcal{Y}$. The model $f$ might take advantage of the backdoor paths to make decisions instead of struggling to discover the real cause-and-effect relationships (Sui et al., 2022; Yang et al., 2021). Our goal is to eliminate the backdoor paths.
3.1.2. Frontdoor Adjustment
In statistics (Murphy, 2012), a simple way to exclude the variable $\mathcal{C}$ in the SCM in Figure 1 is to marginalize it out. However, marginalization requires $\mathcal{C}$ to be observable and measured by the marginal distribution $P(\mathcal{C})$, but in spatiotemporal imputation, $\mathcal{C}$ is usually unknown and difficult to measure. Rather than directly marginalizing $\mathcal{C}$ out, we resort to the frontdoor adjustment (Pearl and Mackenzie, 2018), which uses Pearl’s do-calculus (Pearl and Mackenzie, 2018) to block the backdoor paths. We follow the three steps of the frontdoor adjustment.
  1. (1)
    Remove the backdoor path from $\mathcal{G}$ to $\mathcal{H}$. Given $\mathcal{G}$, there is no backdoor path from $\mathcal{G}$ to $\mathcal{H}$. Note that $\mathcal{G}$ cannot reach $\mathcal{H}$ via $\mathcal{G}\leftarrow\mathcal{C}\rightarrow\mathcal{Y}\leftarrow\mathcal{H}$ according to the d-separation theory (Murphy, 2012). Therefore, we have:
    (3) $P(\mathcal{H}|do(\mathcal{G}))=P(\mathcal{H}|\mathcal{G}).$
  2. (2)
    Remove the backdoor path from $\mathcal{H}$ to $\mathcal{Y}$. There is a backdoor path between $\mathcal{H}$ and $\mathcal{Y}$: $\mathcal{H}\leftarrow\mathcal{G}\leftarrow\mathcal{C}\rightarrow\mathcal{Y}$. This backdoor path can be blocked by marginalizing out $\mathcal{G}$:
    (4) $P(\mathcal{Y}|do(\mathcal{H}))=\sum_{\mathbf{G}}P(\mathcal{Y}|\mathcal{H},\mathbf{G}_{M})P(\mathbf{G}).$
  3. (3)
    Combine the results of the above two steps:
    (5) $\begin{split}P(\mathcal{Y}|do(\mathcal{G}))&=\sum_{i,t}P(\mathbf{h}_{i,j}|do(\mathcal{G}))P(\mathcal{Y}|do(\mathbf{h}_{i,t}))\\ &=\sum_{i,t}P(\mathbf{h}_{i,t}|\mathcal{G})\sum_{\mathbf{G}}P(\mathcal{Y}|\mathbf{h}_{i,t},\mathbf{G})P(\mathbf{G}).\end{split}$
In general, $P(\mathbf{h}_{i,t}|\mathcal{G})$ can be viewed as the encoder $f^{E}$, and the rest part, including $\sum_{\mathbf{G}}P(\mathcal{Y}|\mathbf{h}_{i,t},\mathcal{G})P(\mathbf{G})$ and the sum over all data points $\sum_{i,t}$, can be viewed as the decoder $f^{D}$. In the next subsection, we show how to implemet Equation (5).
3.2. Architecture of Casper
In this subsection, based on Equation (5), we propose a novel Cusality-Aware Spatiotemporal Graph Neural Network (Casper). We first present an overview of Casper  which comprised of a Prompt Based Decoder (PBD) and an encoder with Spatiotemporal Causal Attention (SCA). Next, we elaborate PBD and SCA in detail.
Figure 2. Overview of Casper.
3.2.1. Overview
Figure 2 shows an overview of Casper. The encoder $f^{E}$ is comprised of an input project and $L$ layers of the combination of skip project, transformer, SCA and add & norm. Let $\mathbf{m}\in\mathbb{R}^{d}$ be the embedding for the missing points. The input project module encodes the raw input $\mathbf{G}=(\mathbf{X},\mathbf{A},\mathbf{M})$ into $\mathbf{H}^{(0)}$ via:
(6) $\mathbf{H}^{(0)}=\text{MLP}(\mathbf{X})\odot\mathbf{M}+\mathbf{m}\odot(1-\mathbf{M}),$
where MLP stands for Multi-Layer Perceptron and $\odot$ denotes the Hadamard product. The skip project module prevents gradient vanishing and improves the performance by injecting the $\mathbf{G}$ into the embeddings from the previous layer $\mathbf{H}^{(l-1)}$:
(7) $\mathbf{H}^{skip}=\mathbf{H}^{(l-1)}+\text{MLP}(\mathbf{X})\odot\mathbf{M}+\mathbf{m}\odot(1-\mathbf{M}).$
The transformer encoder layer (Vaswani et al., 2017) learns temporal information for each time series within $\mathbf{H}^{skip}$:
(8) $\mathbf{H}^{in}=\text{Transformer}(\mathbf{H}^{skip}).$
SCA discovers spatiotemporal causal relationships among embeddings based on $\mathbf{A}$, and encodes causal information into embeddings:
(9) $\mathbf{H}^{out}=\text{SCA}(\mathbf{H}^{in},\mathbf{A}).$
The final embeddings of the $l$-th layer is given by:
(10) $\mathbf{H}^{(l)}=\text{LayerNorm}(\mathbf{H}^{in}+\mathbf{H}^{out}).$
Given the embeddings $\mathbf{H}=\mathbf{H}^{(L)}$ obtained by the encoder $f^{E}$, the PBD module in $f^{D}$ generates the predictions $\hat{\mathbf{Y}}=\{\hat{y}_{i,t}\}_{i=1,t=1}^{N,T}$:
(11) $\hat{\mathbf{Y}}=\text{PBD}(\mathbf{H}).$
Figure 3. Prompt Based Decoder (PBD).
3.2.2. Prompt Based Decoder
Suppose we are given the input $\mathbf{G}$ and the target is to recover $\mathcal{Y}=y_{i,t}$. In Equation (5), the decoder $f^{D}$ is comprised of (1) a sum over all possible $\mathbf{G}^{\prime}$ in the dataset $\sum_{\mathbf{G}^{\prime}}P(y_{i,t}|\mathbf{h}_{i^{\prime},t^{\prime}},\mathbf{G}^{\prime})P(\mathbf{G}^{\prime})$ and (2) a sum over all data points $\sum_{i^{\prime},t^{\prime}}$ in $\mathbf{G}$. For (2), since the encoders nowadays, e.g., Transformer (Vaswani et al., 2017), are so powerful that could encode sufficient context information $\mathbf{G}$ in $\mathbf{h}_{i,t}$, and thus the decoder $f^{D}$ could only take $\mathbf{h}_{i,t}$ as input (Marisca et al., 2022), instead of all possible $\mathbf{h}_{i^{\prime},t^{\prime}}$. Therefore, we could drop $\sum_{i^{\prime},^{\prime}t}$ and only implement $\sum_{\mathbf{G}^{\prime}}P(y_{i,t}|\mathbf{h}_{i,t},\mathbf{G}^{\prime})P(\mathbf{G}^{\prime})$.
Now, the challenge is how to implement the sum over $\mathbf{G}^{\prime}$. Simple solutions include randomly sampling a set from the training data, or clustering $\mathbf{G}^{\prime}$ into $K$ clusters and using the cluster centers as an approximation. However, random sampling could be unstable in practice, and clustering requires extra pre-trained models to extract embeddings of $\mathbf{G}^{\prime}$ in advance. Inspired by (Jia et al., 2022), which uses prompts to capture the context information of the downstream task, we introduce a Prompt Based Decoder (PBD) to automatically capture the global context information of the dataset during model training.
An illustration of the proposed PBD is shown in Figure 3. $\mathcal{Q}=\mathbf{h}_{i,t}$ is the query, $\mathcal{P}=\{\mathbf{p}_{n}\}_{n=1}^{N_{P}}$ is the set of learnable prompts. In Figure 3, the Project is a linear function followed with a LayerNorm (Ba et al., 2016). Details of the scaled dot-product attention can be found in (Vaswani et al., 2017), and it can be easily extended into multi-head version as in (Vaswani et al., 2017).
3.2.3. Spatiotemporal Causal Attention
Attention functions (Definition 2.4) have become indispensable in deep learning models (Vaswani et al., 2017), which could effectively capture the context information for the target embedding. Although attention scores could show the correlation between embeddings, yet correlation does not necessarily imply causality and thus sometimes could induce undesired non-causal information into embeddings (Sui et al., 2022). Based on the frontdoor adjustment, PBD could eliminate the impact of the unknown confounders $\mathcal{C}$ and ensure the causality between the input $\mathcal{G}$ and the output $\mathcal{Y}$ by summing over $\mathbf{G}$, $i$ and $t$. However, it guides the attention functions to discover the causal relationships at a high level, and thus the learned causal relationships, i.e., attention scores, might still be dense and a little bit difficult to interpret. To directly guide the model to discover the sparse causal relationships, we first define the causality for embeddings (Definition 3.1) based on the Granger causality (Granger, 1969) (Definition 2.3), and then introduce a novel Spatiotemporal Causal Attention (SCA) module to discover the sparse causality between embeddings.
Definition 0 (Granger Causality for Embeddings).
Denote the target embedding and the set of context embeddings as $\mathcal{Q}=\mathbf{h}_{i,t}^{in}$ and $\mathcal{K}=\{\mathbf{h}_{i^{\prime},t^{\prime}}^{in}\}$. Let $\mathbf{h}_{i,t}^{out}=f_{i,t}(\mathbf{h}_{i,t}^{in};\{\mathbf{h}_{i^{\prime},t^{\prime}}^{in}\})$ be an embedding updating function. If there is a $\mathbf{h}_{j,r}^{in}\in\mathcal{K}$ s.t. changing the value of $\mathbf{h}_{j,r}^{in}$ will change the value of $\mathbf{h}_{i,t}^{out}$, then $\mathbf{h}_{j,r}^{in}$ Granger causes $\mathbf{h}_{i,t}^{out}$.
Let $f_{i,t}$ be an attention function as shown in Definition 2.4:
(12) $\mathbf{h}_{i,t}^{out}=\sum_{i^{\prime},t^{\prime}}\alpha_{i^{\prime},t^{\prime}}\mathbf{v}_{i^{\prime},t^{\prime}},\quad\mathbf{v}_{i^{\prime},t^{\prime}}=f_{v}(\mathbf{h}_{i,t}^{in},\mathbf{h}_{i^{\prime}t^{\prime}}^{in}),$
where $\alpha_{i^{\prime},t^{\prime}}$ is the attention weight, and $f_{v}$ is the message function. According to Definition 3.1, if $\alpha_{j,r}\neq 0$, then $\mathbf{h}_{j,r}^{in}$ Granger causes $\mathbf{h}_{i,t}^{out}$; otherwise if $\alpha_{j,r}=0$, then $\mathbf{h}_{j,r}^{in}$ does not Granger cause $\mathbf{h}_{i,t}^{out}$. In practice, without directly manipulating, $\alpha_{i^{\prime},t^{\prime}}>0$ holds for many noise messages $\mathbf{v}_{i^{\prime},t^{\prime}}$, as shown in Figure 4(c)-4(d) in our experiments.
Figure 4. Spatiotemporal Causal Attention (SCA). Corr. Weight and Causal Prob. correspond to Equation (14)(16).
To further enforce the weights of the noisy points to be zero and discover Granger causality, we propose a novel SCA. As shown in Figure 4, SCA is comprised of two components: (1) a spatiotemporal graph attention function (orange), which learns the correlation between embeddings; (2) a causal gate (blue), which discovers the causal and non-causal relationships.
Let $\mathcal{Q}=\mathbf{h}_{i,t}^{in}$ be the target query embedding, and $\mathcal{N}(i)$ be the neighbors of the $i$-th sensor, i.e., $\mathbf{A}[i,i^{\prime}]\neq 0$. Denote the context keys as $\mathcal{K}=\{\mathbf{h}_{i^{\prime},t^{\prime}}^{in}\}$ and the message values from $\mathcal{N}(i)$ to the point $(i,t)$ as $\mathcal{V}=\{\mathbf{v}_{i^{\prime},t^{\prime}}\}$, where $i^{\prime}\in\mathcal{N}(i)$, $t^{\prime}\in[1,\cdots,T]$ and $\mathbf{v}_{i^{\prime},t^{\prime}}=\text{MLP}([\mathbf{h}_{i,t}^{in};\mathbf{h}_{i^{\prime},t^{\prime}}^{in}])$. We define SCA as:
(13) $\mathbf{h}_{i,t}^{out}=\frac{1}{Z}\sum_{i^{\prime}\in\mathcal{N}(i)}\sum_{t^{\prime}=1}^{T}\beta_{i^{\prime},t^{\prime}}\alpha_{i^{\prime},t^{\prime}}\mathbf{v}_{i^{\prime},t^{\prime}},$
where $Z\in\mathbb{R}$ is a normalization factor, $\alpha_{i^{\prime},t^{\prime}}$ is the correlation weight, $\beta_{i^{\prime},t^{\prime}}\sim\text{Bernoulli}(\rho_{i^{\prime},t^{\prime}})$ is the causal gate, and $\rho_{i^{\prime},t^{\prime}}$ is the probability that $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}$ Granger causes $\mathbf{h}_{i,t}^{out}$. According to Definition 3.1, if $\beta_{i^{\prime},t^{\prime}}\cdot\alpha_{i^{\prime},t^{\prime}}>0$, then $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}$ Granger causes $\mathbf{h}_{i,t}^{out}$; otherwise, there is no Granger causality between $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}$ and $\mathbf{h}_{i,t}^{out}$. The function of calculating correlation weight $\alpha_{i^{\prime},t^{\prime}}$ is given by:
(14) $\alpha_{i^{\prime}t^{\prime}}$ $=\frac{\exp(s(\mathbf{h}_{i,t};\mathbf{h}_{i^{\prime},t^{\prime}}))}{\sum_{j\in\mathcal{N}(i)}\sum_{r=1}^{T}\exp(s(\mathbf{h}_{i,t};\mathbf{h}_{j,r}))},$
(15) $s(\mathbf{h}_{i,t}^{in};\mathbf{h}_{i^{\prime},t^{\prime}}^{in})$ $={(\mathbf{W}_{Q}\mathbf{h}_{i,t}^{in})^{T}(\mathbf{W}_{K}\mathbf{h}_{i^{\prime},t^{\prime}}^{in}})/{\sqrt{d}},$
where $\mathbf{W}_{Q},\mathbf{W}_{K}\in\mathbb{R}^{d\times d}$ are learnable weights, and $d$ is the size of the hidden dimension. We build a neural network to learn the probability $\rho_{i^{\prime},t^{\prime}}$ of the causal gate $\beta_{i^{\prime},t^{\prime}}$:
(16) $\rho_{i^{\prime},t^{\prime}}=\sigma(\mathbf{W}_{c}[\mathbf{W}_{Q_{c}}\mathbf{h}_{i,t}^{in};\mathbf{W}_{K_{c}}\mathbf{h}_{i^{\prime},t^{\prime}}^{in}]),$
where $\mathbf{W}_{c}\in\mathbb{R}^{1\times 2d},\mathbf{W}_{Q_{c}},\mathbf{W}_{K_{c}}\in\mathbb{R}^{d\times d}$ are learnable weights, and $\sigma$ is the Sigmoid activation function.
There are two practical issues of directly using $\rho$ in the above equation. First, the sampling operation $\beta_{i^{\prime},t^{\prime}}\sim\text{Bernoulli}(\rho_{i^{\prime},t^{\prime}})$ is in-differentiable. To address this issue, we use the differentiable re-parameterization technique Gumbel-Softmax (Jang et al., 2016) to obtain $\beta_{i^{\prime},t^{\prime}}$:
(17) $\beta_{i^{\prime},t^{\prime}}=\frac{\exp((\log\rho_{i^{\prime},t^{\prime}}+g)/\tau)}{\exp((\log\rho_{i^{\prime},t^{\prime}}+g)/\tau)+\exp((\log(1-\rho_{i^{\prime},t^{\prime}})+g)/\tau)},$
where $g=-\log(-\log(u))$, $u\sim\text{Uniform}(0,1)$, and $\tau$ is the temperature parameter.
Second, if $\rho_{i^{\prime},t^{\prime}}$ is not close to 0 or 1, the model’s decision could be ambiguous during inference. For example, if $\rho_{i^{\prime},t^{\prime}}=0.2$, then for the same input data, for 20% time, the model shows $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}$ Granger causes $\mathbf{h}_{i,t}^{out}$, and for the other 80% time, the model shows $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}$ does not Granger cause $\mathbf{h}_{i,t}^{out}$. To avoid such an ambiguous situation, we enforce $\rho_{i^{\prime},t^{\prime}}\rightarrow 0/1$ by placing the $l_{1}$ regularization over $\rho_{i^{\prime},t^{\prime}}$.
It can be theoretically proven that $\rho_{i^{\prime},t^{\prime}}$ will converge to 0 or 1 (see Section 3.3.1). Additionally, in practice, the correlation weight $\alpha$ can be easily extended to the multi-head version as in (Vaswani et al., 2017).
3.2.4. Loss Function
Casper is trained by the masked MAE:
(18) $\mathcal{L}=\sum_{i=1,t=1}^{N,T}m_{i,t}\cdot|{y}_{i,t}-\hat{y}_{i,t}|+\lambda||\Phi||_{1},$
where $m_{i,t}$ is the mask, $y_{i,t}$ is the ground-truth value, $\hat{y}_{i,t}$ is the predicted value, $\Phi$ is the set of all $\rho$ in the spatiotemporal causal attention, $\lambda$ is a tunable coefficient, and $||\cdot||_{1}$ denotes the $l_{1}$ norm.
3.3. Framework Analysis
In this subsection, we provide further analysis of the proposed Casper, including theoretical analysis and complexity analysis.
3.3.1. Theoretical Analysis
We theoretically prove that $\rho_{i^{\prime},t^{\prime}}$ in Equation (16) will converge to 0 or 1 in Theorem 1. Therefore, $\rho_{i^{\prime},t^{\prime}}$ indicates the Granger causality. If $\rho_{i^{\prime},t^{\prime}}=0$, then $\beta_{i^{\prime},t^{\prime}}\cdot\alpha_{i^{\prime},t^{\prime}}=0$, showing that $\mathbf{h}^{in}_{i^{\prime},t^{\prime}}$ does not Granger cause $\mathbf{h}^{out}_{i,t}$. Moreover, from the proof of Theorem 1, it can be observed that $\rho_{i^{\prime},t^{\prime}}$ is actually a gradient based explanation (see Remark 1), which determines causal and non-causal relationships based on the gradients. Compared with the classic gradient based explanation methods (Selvaraju et al., 2017), which needs extra steps to calculate gradients after the model is trained, the proposed $\rho$ has two advantages: (1) $\rho$ does not require extra steps for calculating derivatives, and the value of $\rho$ could directly provide the explanation; (2) the parameters associated with $\rho$ are jointly trained with the model, and thus it can guide the model to focus on the most important relationships during training.
Theorem 1 (Convergence of $\rho$).
$\rho$ could converge to 0 or 1 by updating its parameters based on $\mathcal{L}$ in Equation (18).
Proof Sketch.
Due to space limitation, we only provide a proof sketch here. For simplicity, let’s only consider the loss for a single point $(i,t)$, where $m_{i,t}=1$ and ${y}_{i,t}-\hat{y}_{i,t}>0$:
(19) $\mathcal{L}_{i,t}={y}_{i,t}-\hat{y}_{i,t}+\lambda||\Phi||_{1}$
Since Equation (16) is essentially a linear function with a Sigmoid activation, we can rewrite it as:
(20) $\rho_{i^{\prime},t^{\prime}}=\sigma(\mathbf{w}^{T}\mathbf{h}^{in}),\quad\mathbf{h}^{in}=[\mathbf{h}_{i,t}^{in};\mathbf{h}_{i^{\prime},t^{\prime}}^{in}],\quad\mathbf{w}\in\mathbb{R}^{2d}$
Let $w_{j}=\mathbf{w}[j]$ and $h_{j}^{in}=\mathbf{h}^{in}[j]$, then the gradient at $w_{j}$ is:
(21) $\begin{split}\frac{\partial\mathcal{L}_{i,t}}{\partial w_{j}}&=\frac{\partial{y}_{i,t}-\hat{y}_{i,t}}{\partial\mathbf{h}_{i,t}^{out}}\cdot\frac{\partial\mathbf{h}_{i,t}^{out}}{\partial\beta_{i^{\prime},t^{\prime}}}\cdot\frac{\partial\beta_{i^{\prime},t^{\prime}}}{\partial\rho_{i^{\prime},t^{\prime}}}\cdot\frac{\partial\rho_{i^{\prime},t^{\prime}}}{\partial w_{j}}+\lambda\frac{\partial\rho_{i^{\prime},t^{\prime}}}{w_{j}}\\ &=(\lambda-g)\rho_{i^{\prime},t^{\prime}}(1-\rho_{i^{\prime},t^{\prime}})h^{in}_{j}\end{split}$
(22) $\begin{split}g&=\frac{\partial\hat{y}_{i,t}}{\partial\mathbf{h}_{i,t}^{out}}\cdot\frac{\partial\mathbf{h}_{i,t}^{out}}{\partial\beta_{i^{\prime},t^{\prime}}}\cdot\frac{\partial\beta_{i^{\prime},t^{\prime}}}{\partial\rho_{i^{\prime},t^{\prime}}}=\frac{\partial\hat{y}_{i,t}}{\partial\mathbf{h}_{i,t}^{out}}\cdot\frac{\alpha_{i^{\prime},t^{\prime}}}{Z}\mathbf{v}_{i^{\prime},t^{\prime}}\cdot 1\end{split}$
In gradient decent, the updating function of $w_{j}$ is:
(23) $w_{j}^{(k+1)}=w_{j}^{(k)}-\eta(\lambda-g)\rho_{i^{\prime},t^{\prime}}^{(k)}(1-\rho_{i^{\prime},t^{\prime}}^{(k)})h^{in}_{j}$
where $k$ is the iteration index and $\eta>0$ is the learning rate. As we only consider the parameters $\mathbf{w}$ of $\rho$, for simplicity, let’s fix all other parameters in the model. Therefore, ${\partial\hat{y}_{i,t}}/{\partial\mathbf{h}_{i,t}^{out}},\mathbf{v}_{i^{\prime},t^{\prime}},\alpha_{i^{\prime},t^{\prime}}$ are fixed. The normalization factor $Z\leq 1$ and $Z=1$ if and only if $\beta_{i^{\prime},t^{\prime}}=1$ for $\forall i^{\prime},t^{\prime}$. $\rho_{i^{\prime},t^{\prime}}(1-\rho_{i^{\prime},t^{\prime}})>0$ since $\rho_{i^{\prime},t^{\prime}}\in(0,1)$. Now, suppose $g>\lambda$, if $h_{j}^{in}>0$, then ${w}^{(k+1)}_{j}>{w}^{(k)}_{j}$. When $k\rightarrow\infty$, $w_{j}^{(k)}\rightarrow+\infty$. Otherwise, if $h_{j}^{in}<0$, then ${w}^{(k+1)}_{j}<{w}^{(k)}_{j}$. When $k\rightarrow\infty$, $w_{j}^{(k)}\rightarrow-\infty$. Therefore, when $k\rightarrow\infty$, $\mathbf{w}^{T}\mathbf{h}\rightarrow+\infty$. As a result, $\rho_{i^{\prime},t^{\prime}}^{(k)}=\sigma({\mathbf{w}^{(k)}}^{T}\mathbf{h}^{in})\rightarrow 1$, as $k\rightarrow\infty$. Similarly, if $g<\lambda$, then $\rho_{i^{\prime},t^{\prime}}^{(k)}$ will converge to 00. ∎
Remark 1 ($\rho$ is a gradient based explanation).
From the proof of Theorem 1, it can be noted that, if $g>\lambda$, then $\rho_{i^{\prime},t^{\prime}}\rightarrow 1$; if $g<\lambda$, then $\rho_{i^{\prime},t^{\prime}}\rightarrow 0$. This phenomenon reflects that $\rho_{i^{\prime},t^{\prime}}$ serves as a binary indicator showing whether the gradient at $\rho_{i^{\prime},t^{\prime}}$ is greater or less than the threshold $\lambda$.
3.3.2. Complexity Analysis
The complexities of the input project, the skip project and the transformer layer are $O(NT)$, $O(NT)$ and $O(NT^{2})$. The complexities of SCA and PBD are $O(ET)$ and $O(NN_{P}T)$, where $E$ is the number of edges in the sensor network, $N_{P}$ is the number of prompts. The overall complexity is $O(\max(E,NT,NN_{P})T)$. In SCA, if we calculate attention weights and causal gates for each pair of the data point without using $\mathbf{A}$, then the complexity will be extremely high: $O(N^{2}T^{2})\gg O(ET)$.
4. Experiments
Table 1. Performance (MAE, MSE) of different methods.
General Missing Point Missing Block Missing
AQI-36 AQI METR-LA PEMS-BAY METR-LA PEMS-BAY
MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
Mean 53.48$\pm$0.00 4578.08$\pm$00.00 39.60$\pm$0.00 3231.04$\pm$00.00 7.56$\pm$0.00 142.22$\pm$0.00 5.42$\pm$0.00 86.59$\pm$0.00 7.48$\pm$0.00 139.54$\pm$0.00 5.46$\pm$0.00 87.56$\pm$0.00
KNN 30.21$\pm$0.00 2892.31$\pm$00.00 34.10$\pm$0.00 3471.14$\pm$00.00 7.88$\pm$0.00 129.29$\pm$0.00 4.30$\pm$0.00 49.80$\pm$0.00 7.79$\pm$0.00 124.61$\pm$0.00 4.30$\pm$0.00 49.90$\pm$0.00
MF 30.54$\pm$0.26 2763.06$\pm$63.35 26.74$\pm$0.24 2021.44$\pm$27.98 5.56$\pm$0.03 113.46$\pm$1.08 3.29$\pm$0.01 51.39$\pm$0.64 5.46$\pm$0.02 109.61$\pm$0.78 3.28$\pm$0.01 50.14$\pm$0.13
MICE 30.37$\pm$0.09 2594.06$\pm$07.17 26.98$\pm$0.10 1930.92$\pm$10.08 4.42$\pm$0.07 55.07$\pm$1.46 3.09$\pm$0.02 31.43$\pm$0.41 4.22$\pm$0.05 51.07$\pm$1.25 2.94$\pm$0.02 28.28$\pm$0.37
VAR 15.64$\pm$0.08 833.46$\pm$13.85 22.95$\pm$0.30 1402.84$\pm$52.63 2.69$\pm$0.00 21.10$\pm$0.02 1.30$\pm$0.00 6.52$\pm$0.01 3.11$\pm$0.08 28.00$\pm$0.76 2.09$\pm$0.10 16.06$\pm$0.73
rGAIN 15.37$\pm$0.26 641.92$\pm$33.89 21.78$\pm$0.50 1274.93$\pm$60.28 2.83$\pm$0.01 20.03$\pm$0.09 1.88$\pm$0.02 10.37$\pm$0.20 2.90$\pm$0.01 21.67$\pm$0.15 2.18$\pm$0.01 13.96$\pm$0.20
BRITS 14.50$\pm$0.35 662.36$\pm$65.16 20.21$\pm$0.22 1157.89$\pm$25.66 2.34$\pm$0.00 16.46$\pm$0.05 1.47$\pm$0.00 7.94$\pm$0.03 2.34$\pm$0.01 17.00$\pm$0.14 1.70$\pm$0.01 10.50$\pm$0.07
ST-Transformer 11.98$\pm$0.53 557.22$\pm$46.52 18.11$\pm$0.25 1135.46$\pm$89.27 2.16$\pm$0.00 13.66$\pm$0.03 0.74$\pm$0.00 1.96$\pm$0.03 3.54$\pm$0.00 52.22$\pm$0.99 1.70$\pm$0.02 20.37$\pm$0.43
GRIN 12.08$\pm$0.47 523.14$\pm$57.17 14.73$\pm$0.15 775.91$\pm$28.49 1.91$\pm$0.00 10.41$\pm$0.03 0.67$\pm$0.00 1.55$\pm$0.01 2.03$\pm$0.00 13.26$\pm$0.05 1.14$\pm$0.01 6.60$\pm$0.10
SPIN 11.77$\pm$0.54 455.53$\pm$12.27 13.92$\pm$0.15 773.60$\pm$26.64 1.90$\pm$0.01 18.47$\pm$0.31 0.70$\pm$0.01 1.91$\pm$0.01 1.98$\pm$0.01 18.47$\pm$0.31 1.06$\pm$0.02 7.42$\pm$0.16
Casper 10.09$\pm$0.13 396.16$\pm$12.94 13.30$\pm$0.06 658.07$\pm$4.88 1.84$\pm$0.00 9.99$\pm$0.01 0.65$\pm$0.00 1.63$\pm$0.01 1.92$\pm$0.01 11.98$\pm$0.23 1.00$\pm$0.00 5.37$\pm$0.04
4.1. Experimental Setup
Datasets. Three public real-world datasets are used to evaluate Casper. AQI (Zheng et al., 2015) is an hourly record of air pollutants from 437 air quality monitoring stations in China from May 2014 to April 2015. Besides, we also use the popular AQI-36 (Cao et al., 2018) which is a reduced version of AQI containing records from 36 sensor stations scatter around Beijing. METR-LA (Li et al., 2018) contains traffic speed time series collected from 207 sensors on highways in Los Angeles for 4 months. PEMS-BAY (Li et al., 2018) is traffic speed time series collected from 325 sensors on highways in San Francisco Bay Area for 6 months. The records in both METR-LA and PEMS-BAY are collected every 5 minutes. For AQI, METR-LA and PEMS-BAY, the temporal window is set as $T=24$, and for AQI-36, the temporal window is 36. Following (Marisca et al., 2022), the sensor network is built based on the geographical distances between sensor stations.
For AQI and AQI-36, we use the evaluation masks in (Yi et al., 2016; Marisca et al., 2022) which simulates the real missing data distribution in the datasets. We refer to this setting as the general missing. For METR-LA and PEMS-BAY, we consider both point missing and block missing settings as in (Marisca et al., 2022). In point missing, 25% data points are masked out. In block missing, 5% spatial blocks and 0.15% temporal blocks ranging from 1 hour to 4 hours are masked out.
Evaluation Metrics. We use the standard Mean Absolute Error (MAE) and Mean Squared Error (MSE) between the ground truth values and the imputed values as the evaluation metrics.
Baselines. We consider three groups of baselines. (1) Traditional statistical methods: the mean value of the sequence (MEAN); neighbor mean (KNN); matrix factorization (MF); multiple imputation using chained equations (MICE (White et al., 2011)); vector auto-regression (VAR); (2) Early deep learning models: rGAIN (Cini et al., 2021): an adversarial method similar to (Miao et al., 2021; Luo et al., 2018); BRITS (Cao et al., 2018): a bidirectional RNN method; (3) Recent deep learning models: ST-Transformer (Marisca et al., 2022): a spatiotemporal extension of the original Transformer (Vaswani et al., 2017); GRIN (Cini et al., 2021): a graph enhanced recurrent neural network; SPIN (Marisca et al., 2022): a spatiotemporal graph attention based imputation model. Whenever possible, the results of baselines are copied from the corresponding paper.
Implementation Details. We set the size of embeddings as 32. The numbers of layers for the encoder for AQI-36 and other datasets are 2 and 4 respectively. Following (Marisca et al., 2022), we use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.0008 and a cosine scheduler to train the model. The maximum number of epochs is 300 and the patience of early stopping is 40 epochs. Batch size is fixed as 8. During training, $p\in[0.2,0.5,0.8]$ data points are randomly masked out for each batch, and the loss is calculated based on these masked points. Our code is implemented based on PyG (Fey and Lenssen, 2019) and Torch Spatiotemporal (Cini and Marisca, 2022). Code will be released upon publication.
4.2. Overall Performance
The overall performance (MAE and MSE) of different methods is presented in Table 1. The upper, middle and lower groups of baselines are the traditional statistical methods, the RNN methods and the recent methods (Transformer based and graph based methods). Generally speaking, the RNN methods perform better than the statistical methods, and the recent methods further outperform the RNN methods. When imputing a data point, these methods exploit all the available information in the context, without identifying the causal relationships between the data point and the context. However, it is inevitable that some confounders are included in the data, such as noise and non-causal shortcut edges. Over-reliance on the confounders could lead to overfitting and make the model susceptible to noise. The proposed Casper could effectively remove the impact of confounders. As shown in Table 1, Casper achieves the lowest overall MAE and MSE scores and also has lower standard deviations, demonstrating the effectiveness of enforcing the model to discover causality during training.
Table 2. Ablation study on the AQI-36 dataset.
MAE MSE
Casper 10.09$\pm$0.13 396.16$\pm$12.94
[…]
w/o SCA (i.e., w/o $\beta$) 10.45$\pm$0.17 426.64$\pm$34.41
w/o PBD, SCA 10.84$\pm$0.20 472.24$\pm$42.77
w/o PBD, SCA, $\mathbf{A}$ 14.86$\pm$0.21 767.96$\pm$25.32
Prompts $\rightarrow$ K-means Centers (max) 10.18$\pm$0.14 427.20$\pm$28.84
Prompts $\rightarrow$ K-means Centers (avg) 10.23$\pm$0.15 421.43$\pm$19.40
Prompts $\rightarrow$ Sampling (max) 10.25$\pm$0.18 421.78$\pm$33.26
Prompts $\rightarrow$ Sampling (avg) 10.56$\pm$0.19 474.27$\pm$50.55
(a) Input
(b) Casper
(c) Casper w/o SCA
(d) Casper w/o SCA, PBD
(e) Input
(f) Casper
(g) Casper w/o SCA
(h) Casper w/o SCA, PBD
Figure 5. Input matrices and the associated attention maps of Casper and its ablated versions. Diamonds are the query points.
4.3. Ablation Study
In this subsection, we study the impact of different components in Casper, and the results are shown in Table 2.
Effectiveness of PBD,SCA and $\mathbf{A}$. In the upper part of Table 2, we investigate the impact of the Prompt-Based Decoder (PBD), the Spatiotemporal Causal Attention (SCA) and the sensor network $\mathbf{A}$. First, significant performance drop on MAE and MSE can be observed when we remove PBD (replace PBD with a MLP) and/or SCA (remove the causal gate $\beta$) from the full model Casper, indicating the effectiveness of the frontdoor adjustment and the causal gate for improving the overall performance. It is also worth noting that, on MSE, the standard deviations of the ablated versions of Casper (w/o PBD and/or SCA) are significantly higher than the full model Casper, demonstrating that enforcing the model to focus on the causality could improve its robustness. Second, although $\mathbf{A}$ contains non-causal relationships among sensors, it still has critical contributions to the overall performance, demonstrating the necessity of considering $\mathbf{A}$ when imputing missing values.
Effectiveness of the Prompts. In the lower part of Table 2, we demonstrate the effectiveness of using the learnable prompts to capture the global contextual information of the datasets by replacing the prompts with other approximations, including K-means cluster centers and randomly sampled data. To obtain the other two approximations, we first pre-train an imputation model, i.e., Casper without SCA and PBD, and then extract the embedding of each training sample $\mathbf{G}$ by applying average/max pooling over the embeddings of all the points in $\mathbf{G}$. For the cluster center approximation, we apply K-means over the emebddings to obtain 1,000 cluster centers. For the sampling approximation, we randomly sample 1,000 embeddings for each training sample. Compared with the prompts, the cluster centers and randomly sampled embeddings not only perform worse but also require extra efforts to obtain embeddings, which demonstrates the superiority of prompts.
(a) Causal relations in Figure 4(b).
(b) Causal relations in Figure 4(f).
Figure 6. Discovered causal relationships.
4.4. Visualization
In this subsection, we provide visualization to further analyze Casper’s ability for discovering the causality.
[…]
Discovered Causal Relationships. In Figure 6, we visualize the discovered causal relationships corresponding to Figure 4(b) and 4(f). The AQI-36 dataset is collected from 36 air quality monitoring stations in Beijing, which records PM2.5 levels. We draw the sensors on the map of Beijing according to their latitudes and longitudes. The widths of the arrows correspond to the values of attention scores. It is evident that for the query sensors 14 and 1 in the two figures, not every their nearby sensor has a causal relationship with them. In Figure 5(a), although the sensor 6, 5, 15 are very close to the query 14, Casper discovers that there is no causal relationship among them. We conjecture that this is because the wind blew generally westward and northward in Beijing during the time period corresponding to the input data. In Figure 5(b), the neighboring sensors 24, 15, 6, 20, 9, 13 are not regarded as the causes for the query sensor 1. According to the spatial relationships, the information of the sensors 15, 6, 20, 9, 13 might be included in the causal sensors such as 2, 5 14, 8, since the causal sensors are in between with the query and theses non-causal sensors. However, there is no other sensor between 1 and 24. By taking a closer look at the upper-left corner of the map, we find that the sensors 1 and 24 are separated by the Fragrant Hills. Therefore, the air quality at the sensor 1 might not be directly influenced by the sensor 24 for that the time period corresponding to the input data. The two examples in Figure 6 show that the proposed Casper could effectively discover the causal relationships among sensors, which provides better insights for further data analysis. Additionally, the two examples also show that the simple distance based sensor network construction is biased, which contains many non-causal correlations, since it ignores other factors in the real world, such as wind direction and terrain.
(a) The weight of $l_{1}$ norm $\lambda$.
(b) The number of prompts $N_{P}$.
Figure 7. Sensitivity experiments.
4.5. Sensitivity Analysis
We present the results of sensitivity experiments of $\lambda$ and the number of prompts $N_{P}$ in Figure 6(a)-6(b). For $\lambda$, the lowest MAE can be obtained when $\lambda$ is around 0.001. For $N_{P}$, the best results can be obtained when $N_{P}\in[200,1400]$.
5. Related Work
[…]
Causal Inference. Causality theory (Pearl and Mackenzie, 2018) provides theoretical guidance to design causality aware models. It has been widely explored in the computer vision domain to discover causal relationships (Bengio et al., 2019), generate counter-factual samples (Abbasnejad et al., 2020; Yue et al., 2021; Kocaoglu et al., 2018) and reduce bias (Hu et al., 2021; Yang et al., 2021; Qi et al., 2020). In the graph mining domain, CGI (Feng et al., 2021) studies how to select trustworthy neighbors during inference; CLEAR (Ma et al., 2022a) explores how to generate counterfactual explanations for graph-level prediction models based on Independent Component Analysis (ICA) (Khemakhem et al., 2020); HyperSCI (Ma et al., 2022b) explores Individual Treatment Effect (ITE) on hyper-graphs. NEAT (Ma et al., 2023) investigates the impact of MRSA infection via the Neyman-Rubin potential outcome framework (Rubin, 2005). CAL (Sui et al., 2022) introduces a causal attention learning framework for graph neural networks based on the backdoor adjustment (Pearl and Mackenzie, 2018). There are two differences between Casper and the above graph methods: (1) our setting is the dynamic spatio-temporal setting but their setting is static graph; (2) Casper is based on the frontdoor adjustment and Granger causality, which is fundamentally different from their theoretical basis of causality. In the time series domain, the Granger causality (Granger, 1969) is widely used for analyzing the causality between time series in the forecasting setting. GrID-Net (Wu et al., 2021) leverages the Granger causality to infer regulatory locus–gene links. cLSTM (Tank et al., 2021) and economy-SRU (Khanna and Tan, 2019) integrates the Granger causality with LSTM (Hochreiter and Schmidhuber, 1997) and SRU (Oliva et al., 2017). However, thses methods require the input time series to be fully observed. CUTS (Cheng et al., 2022) is a recently proposed two-stage model, which first imputes missing data and then discovers causality between time series. There are three major differences between Casper and CUTS. (1) CUTS does not consider confounders, while Casper removes confounders via the frontdoor adjustment. (2) CUTS has to re-train the causal model for each input snippet, but Casper does not have such a requirement. (3) CUTS is a two stage model, while Casper is an end-to-end model.
6. Conclusion
In this paper, we review the spatiotemporal time series imputation task via the Structure Causal Model (SCM), which shows the causal relationships among the input, output, embeddings and confounders. The confounders could open shortcut backdoor paths between the input and output, which could mislead the model to learn the non-causal relationships. We use the frontdoor adjustment to block the backdoor paths. Based on the results of the frontdoor adjustment, we proposed a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper), which is comprised of a Prompt Based Decoder (PBD) and an encoder equipped with Spatiotemporal Causal Attention (SCA). The proposed PBD could reduce the impact of the confounders at a high level. The proposed SCA could further discover the sparse causal relationships among embeddings. Theoretical analysis shows that SCA decides causal and non-casual relationships based on the values of gradients. Experimental results on three real-world benchmark datasets show that Casper could outperform the baselines methods, and could also effectively discover the causal relationships.